# photon/.style=decorate, decoration=snake, draw=black

### v1

@@ Abstract
Extraction tests of Compton Form Factors are performed using pseudodata based on experimental data from Deeply Virtual Compton Scattering experiments conducted at Jefferson Lab. The standard Belitsky, Kirchner, and Muller formalism at twist-two is employed, along with a fitting procedure designed to reduce model dependency similar to traditional local fits. The extraction of the Compton Form Factors is performed using both Classical Deep Neural Networks (CDNNs) and Quantum Deep Neural Networks (QDNNs). Comparative studies reveal that QDNNs outperform CDNNs for this application, demonstrating improved predictive accuracy and precision even for limited model complexity. The results demonstrate the potential of QDNNs for future studies in which quantum algorithms can be fully optimized.
@@ I Introduction
Unraveling the internal structure of the nucleon has been a longstanding goal in hadronic physics, with deeply virtual exclusive processes providing a crucial window into its partonic composition. Among these, Deeply Virtual Compton Scattering (DVCS) stands out as a uniquely sensitive probe of Generalized Parton Distributions (GPDs), which encode the spatial and momentum distributions of quarks and gluons. By linking longitudinal momentum fractions to transverse spatial coordinates, GPDs provide critical insights into the orbital angular momentum carried by partons, bridging the gap between experimental observables and the deeper theoretical understanding of nucleon structure.
@@
Despite their fundamental importance, GPDs cannot be directly extracted from experimental measurements due to their integral appearance in DVCS amplitudes. Instead, Compton Form Factors (CFFs) emerge as a more experimentally accessible alternative, encapsulating the convolution of GPDs with perturbative coefficient functions. These complex-valued functions serve as indirect yet powerful observables, enabling the study of quark and gluon distributions through measurable quantities such as cross-sections and asymmetries. However, the extraction of CFFs remains a formidable challenge, as they are embedded in helicity amplitudes and entangled with quantum interference effects. This may necessitate the advancement of analysis techniques to go beyond classical methods, leveraging approaches capable of capturing the quantum correlations inherent in the DVCS process. In this context, quantum deep neural networks (QDNNs) offer a promising approach to the systematic extraction of quantum information pertaining to partonic degrees of freedom, enhancing our ability to measure and interpret the quantum structure of the nucleon.
@@
QDNNs are neural network architectures that take advantage of quantum mechanics to process information. Unlike classical deep learning models that rely on traditional computation, QDNNs employ quantum circuits as trainable components, allowing them to explore complex, high-dimensional function spaces more efficiently. Libraries such as PennyLane and Quantum TensorFlow provide frameworks to construct and train QDNNs using hybrid quantum-classical approaches. These frameworks enable the encoding of classical data into quantum states, manipulation through quantum circuits, and extraction of results via measurement processes. PennyLane is designed to integrate quantum computing with machine learning and automatic differentiation techniques, allowing users to build quantum nodes (or “QNodes”) that can be optimized in conjunction with classical deep learning architectures. Quantum TensorFlow extends Google’s TensorFlow to incorporate quantum computation, enabling quantum variational circuits to be trained using classical gradient descent methods. These tools provide a seamless way to incorporate quantum algorithms into machine learning tasks, facilitating the study of quantum-enhanced feature extraction, pattern recognition, and regression tasks, particularly those of a quantum nature.
@@
One of the compelling applications of QDNNs is in the extraction of kinematically sensitive parameters from helicity amplitudes, such as the CFFs in deeply virtual processes. In DVCS, the helicity amplitudes of the interaction encode critical information about the partonic structure of the nucleon. These amplitudes are complex-valued functions of kinematic variables like the momentum transfer squared, the Bjorken scaling variable, and the photon virtuality. Extracting these CFFs from experimental measurements is challenging because of limitations in constraints, both measured and mathematical, but also because of the inherent interference with the Bethe-Heitler process. The interference term in DVCS and Bethe-Heitler processes emerges because the final state (electron, nucleon, and photon) is produced by two coherent quantum amplitudes that interfere constructively or destructively. This is a direct consequence of quantum superposition and phase coherence. Traditionally, machine learning methods, particularly feed-forward deep neural networks, have been employed to fit experimental data and extract the underlying parameters. However, classical models may not be optimal when dealing with problems involving interference effects, quantum entanglement, and correlations between partonic degrees of freedom. QDNNs provide an alternative approach that naturally accommodates such quantum mechanical correlations. By encoding kinematic inputs into quantum states and using quantum variational circuits to process them, QDNNs can potentially uncover structures in the data that classical networks fail to capture.
@@
By leveraging quantum entanglement and interference, QDNNs can capture nontrivial correlations between observables that are difficult for classical models to learn efficiently. In CFF extraction, classical neural networks have to approximate real and imaginary parts separately, often missing interference effects that are intrinsic to the DVCS amplitude. A QDNN, however, can process this information holistically by leveraging superposition and quantum state evolution, which naturally preserve these quantum correlations. Additionally, quantum-enhanced feature spaces allow for a more expressive representation of the functional dependencies of helicity amplitudes. This is particularly useful in scenarios where experimental uncertainties and systematic errors obscure direct parameter extraction. Since QDNNs operate in a higher-dimensional Hilbert space, they can encode more complex functional mappings, potentially leading to more accurate determinations of observables like beam-spin and target-spin asymmetries, which are crucial in probing the nucleon structure.
With the rapid advancement of near-term quantum processors, such as those developed by IBM, Google, and Rigetti, experimental physics stands at the threshold of utilizing QDNNs for real-world data analysis. A promising approach involves integrating quantum machine learning with Monte Carlo sampling methods to systematically propagate experimental errors through the quantum model, improving the robustness of extracted parameters like CFFs. The application of Quantum DNNs in libraries like PennyLane and Quantum TensorFlow offers a new paradigm for extracting quantum-mechanical observables in high-energy physics. By exploiting quantum feature spaces, entanglement, and parameterized quantum circuits, these models can potentially outperform classical approaches in unraveling the helicity structure of deeply virtual processes like DVCS. As quantum hardware continues to improve, QDNNs may soon become indispensable tools for analyzing quantum field-theoretic processes, leading to deeper insights into the fundamental structure of matter. This approach may enhance the extraction fidelity of critical observables and open new avenues for studying the quantum correlations embedded within the hadronic structure.
@@
In this work, we demonstrate some very basic applications of QDNNs for information extraction and develop some metrics for comparison. Then we perform some extraction testing using a standard local fit involves fitting a single observables at fixed kinematics, across the independent variable $\phi$, the azimuthal angle between the lepton and hadron scattering planes. Local fitting relies on the chosen framework of general approximation, such as twist truncation, leading-order contributions, and the dominance of some CFFs. At the highest approximation, the parametrization of observables, such as the cross-section, in terms of CFFs is not unique as it is convention-dependent since it violates the required symmetries which are restored by contributions that are formally suppressed by powers of $1/Q$, dubbed higher-twist corrections. In this regard, the extracted CFF depends on the choice of the observable formalism [1, 2, 3, 4, 5, 6].
@@
While local fits offer the advantage of being potentially model-independent, extracting unique CFFs for fixed kinematics remains a formidable challenge, particularly when the number of available observables is limited, reducing the constraints in the fit. Since CFFs can vary freely within a high-dimensional parameter space, standard $\chi^{2}$ minimization methods often suffer from inconsistent local minima, leading to significant uncertainties in their extraction, even in the most rigorous analyses. As experimental datasets continue to grow in both quantity and precision, particularly with the inclusion of multiple polarized observables, the issue of data sparsity is gradually being mitigated. However, to fully capitalize on these advancements, it is equally crucial to refine and enhance the extraction techniques, ensuring they are capable of handling the increasing complexity and richness of the data.
@@
We employ a straightforward approach using simple feedforward architectures and, similar to previous work  [7, 8], incorporate the helicity amplitude formalism directly into the loss function during the backpropagation of error. To account for experimental uncertainties, we propagate errors through a Monte Carlo ensemble of replicas [9], where the cross-section data is resampled within its experimental uncertainty. This ensemble generates a distribution of extracted CFFs, with the width of the distribution reflecting both the inherent experimental errors and the uncertainties arising from the fitting procedure.
@@
We focus on the least constrained observable case: the unpolarized DVCS cross-section. We adopt the formalism developed by Belitsky and Müller, specifically the BKM10 framework  [3], which defines the structure of the helicity amplitudes used in the extraction. While this approach is tailored to the BKM10 formalism, the extraction method itself is general and can be adapted to alternative theoretical frameworks as needed. Future work will explore an expanded approach aimed at minimizing potential biases introduced by the choice of helicity amplitude formalism. A brief overview of the BKM10 formalism will be provided in the next section.
@@
In this study, we compare the performance of CDNNs and QDNNs in the extraction of CFFs, using a consistent theoretical framework and a new quantum metric. To ensure a direct comparison, we adopt the BKM10 formalism  [3] for both approaches, maintaining uniformity in the helicity amplitude structure used in the loss function. We also focus on quantifying and matching model complexity. By applying identical pseudo input data and theoretical constraints, we systematically evaluate how QDNNs enhance the extraction process relative to CDNNs, particularly in capturing quantum correlations and mitigating fit ambiguities. The BKM10 formalism, which will be briefly discussed in the next section, provides a well-established foundation for this comparative study, allowing us to isolate the advantages introduced by quantum machine learning techniques.
The remainder of this paper is organized in the following way. In Section II, we provide a detailed description of the theoretical framework for the DVCS channel and describe how CFFs are relevant to the extraction process. In Section VII we present the experimental data to be used to generate the pseudodata. In Section IX, we describe how the pseudodata is produced and perform a standard DNN local fit for comparison. We then perform some basic CDNN and QDNN comparison with regression and classification task. And the perform the extraction of the DVCS CFFs with both CDNN and QDNN. In Section X, we present our conclusions and discuss future outlooks for this research.
@@ II Theory Framework
The DVCS process studies the virtual Compton scattering in deep-inelastic kinematics at the leading order in perturbative QCD. DVCS is dominated by single-quark scattering, and therefore the amplitude can be expressed in terms of the off-forward parton distributions. In the helicity-independent case, a virtual photon scatters from an unpolarized electron beam of energy $k$ off an unpolarized proton leading to the 4-fold photon electroproduction differential cross-section:
@@
The phase space of this process is parameterized by the Bjorken variable, $x_{B}=Q^{2}/2pq$, the four-momentum $q=k-k^{\prime}$ carried by the virtual photon of mass $Q^{2}={-q}^{2}=-(k-k^{\prime})^{2}$, the squared momentum transfer between the initial and final protons $t=\Delta^{2}$ with $\Delta=p^{\prime}-p$ and the lepton energy loss $y=(p\cdot q)/(p\cdot k)$. The azimuthal angle $\phi$ between the leptonic and hadronic planes is defined in the Trento convention [10]. In the cross-section, $\alpha=e^{2}/(4\pi)$ is the fine structure constant, and $\epsilon=2x_{B}M/Q^{2}$ where $M$ is the proton mass.
The photon electroproduction cross-section is sensitive to the coherent interference of the DVCS amplitude with the Bethe-Heitler (BH) amplitude. The DVCS process, namely, the electroproduction of a photon off a proton with the exchange of a photon with virtuality, $Q^{2}$, much larger than the momentum transfer squared $t$, hypothetically provides one of the cleanest probes of the GPDs. The BH is an easily calculated electromagnetic process, where a real photon is emitted by the initial or the final lepton. Therefore both, DVCS and BH channels contain the same final states, which can not be distinguished experimentally, and the total photon electroproduction amplitude squared is given by the linear superposition of these two processes:
@@
The DVCS amplitude encodes the partonic structure of the nucleon given by the Compton tensor $T^{\mu\nu}$  [12] where the observables for the extraction of GPDs emerge. The proven QCD factorization theorem on DVCS [13, 14], permits to express the Compton tensor into perturbative coefficients and the long-distance dynamics that give rise to the GPDs and introduce the handbag diagrams twist expansion. The proven QCD factorization theorem on DVCS [13, 14], which is valid only at values of $t$ such that $-t\ll Q^{2}$, permits separating the Compton tensor into the hard scattering process at short distances between the virtual photon, electron, and quarks, which can be calculated using perturbative QCD, from the soft, long-distance dynamics that give rise to the non-perturbative GPDs that describe the internal structure of hadrons. The soft interactions of the struck parton by the virtual photon with the rest of the proton can be viewed as the exchange of partons of virtuality of the order of $Q^{2}$. Powers of $1/Q$ suppress their contribution to the cross-section of the process as the number of exchanged partons increases. The hard part of scattering processes is described using a perturbative expansion in increasing orders of the strong coupling constant $\alpha_{s}$. This perturbative expression of the hard part in the scattering process is known as the coefficient function.
At leading order (LO) in perturbative QCD and leading twist (LT) approximation, four independent proton GPDs can be accessed in the DVCS process: H, E, $\widetilde{\text{H}}$ and $\widetilde{\text{E}}$. These four GPDs depend on three variables: $x$, $\xi$ and $t$ where the latter two are experimental parameters with the skewness $\xi=x_{B}/(2-x_{B})$ at leading twist and $x\pm\xi$ the light-cone longitudinal momentum fraction of the struck quark before (“$+$” sign) and after (“--” sign) the scattering, such that the difference in the momentum fraction of the quark before absorbing the virtual photon and after the emission of the real photon is given by $2\xi$, resulting in a squared four-momentum transfer $t$ to the nucleon. This gives rise to the division of the GPDs into two distinct regions, each carrying entirely different physical interpretations. In the region where the momentum fraction $x$ exceeds $\xi$ or falls below $-\xi$ the GPDs are interpreted as the amplitudes governing the emission and reabsorption of a parton. Conversely, within region encompassing $-\xi<x<\xi$, GPDs resemble distribution amplitudes governing the emission or absorption of a parton-antiparton pair.
@@
$e_{q}$ is the charge of quarks in units of proton charge and the summation runs over all quark flavors $q$. Using the residue theorem, it is possible to decompose the integral (5) so that each CFF contains two real quantities:
@@
$\Re e\mathcal{F}\left(\xi,t\right)=\sum_{q}\mathcal{P}\int_{0}^{1}\mathrm{~{}d}x\left(\frac{1}{\xi-x}\mp\frac{1}{\xi+x}\right)F_{q}^{[\mp]}\left(x,\xi,t\right),$ (8)
@@
The top sign in equations (5)-(10), corresponds to the unpolarized GPDs ($H,E$) and the bottom sign applies to the polarized GPDs ($\widetilde{H},\widetilde{E}$). Equations (8) and (9) show that observables sensitive to the imaginary part of the CFFs will only contain information along the line $x=\pm\xi$, whereas the real part probes GPD integrals over the momentum fraction $x$. Given the complicated kinematic dependence of the GPDs, retrieving them from CFFs is therefore a major challenge of the field, known as the deconvolution problem [16]. Thus, the maximum model-independent information which can be extracted from the DVCS reaction at leading twist are 8 CFFs, which depend on two variables, $\xi$ and $t$, at QCD leading order. There is an additional $Q^{2}$ dependence in the CFFs (and in the GPDs) if QCD evolution is taken into account.
The unpolarized beam and unpolarized target cross-sections rely on the Belitsky-Kirchner-Muller DVCS formulation (BKM10) [3] and we focus on the case where only twist-2 CFFs enter the cross-section. In this case, the squared DVCS amplitude is a bilinear combination of the CFFs contained in the coefficient $\mathcal{C}^{DVCS}(\mathcal{F},\mathcal{F}^{*})$:
@@
The double helicity-flip gluonic CFFs amplitudes, $C^{\mathcal{I},n}_{-+}(\mathcal{F_{T}})$, are formally suppressed by $\alpha_{s}$ and will be neglected here. $C^{\mathcal{I},n}_{++}(\mathcal{F})$ and $C^{\mathcal{I},n}_{0+}(\mathcal{F}_{eff})$ are the helicity-conserving and the helicity-changing amplitudes respectively. The helicity-changing amplitudes also contain the contribution from twist-3 CFFs implicit in the effective CFFs ($\mathcal{F}_{eff}$). Thus, we chose a framework where only the helicity-conserving amplitudes are considered which contain only pure twist-2 CFFs contributions, similarly to [17]. In that way, we set more constraints on the physics with the same amount of data, giving up to some information at the expense of greater accuracy in the remaining parameters. The helicity-conserving amplitudes are given by:
@@
In this specific scenario, the interference cross-section becomes reliant solely on three real-valued CFFs: $\Re e\mathcal{H}$, $\Re e\mathcal{E}$ and $\Re e\widetilde{\mathcal{H}}$. The real parts of the CFFs $\Re e\mathcal{H}$, $\Re e\mathcal{E}$ and $\Re e\widetilde{\mathcal{H}}$ arise from the dispersive integrals of the corresponding GPDs $H$, $E$, and $\tilde{H}$, representing their convolution with the hard scattering kernel. $\Re e\mathcal{H}$ dominates the DVCS amplitude for unpolarized targets and reflects the spatial-momentum distribution of unpolarized quarks. $\Re e\mathcal{E}$ is sensitive to nucleon helicity-flip and provides access to parton orbital angular momentum via Ji’s sum rule. $\Re e\widetilde{\mathcal{H}}$ encodes the distribution of polarized quarks and governs the longitudinal spin structure, playing a leading role for polarized targets. Additionally, accounting for the constant nature of the pure DVCS cross-section in terms of the azimuthal angle, the total number of parameters that need to be extracted is reduced to four. We chose this framework to make the comparison of CDNN and QDNN extractions more direct. The reduction in the parameter count simplifies the extraction process relative to the full eight-parameter scenario, while still preserving the most relevant aspects of the phenomenological information. This simplified parameter space facilitates a more manageable data analysis and modeling, making a more straightforward demonstration of the extraction steps of the essential parameters from experimental data. A separate future paper will investigate directly the quantum advantage in extracting the complex valued CFFs.

### v2

# Compton Form Factor Extraction using Quantum Deep Neural Networks

@@ Abstract
We present an extraction of Compton Form Factors (CFFs) from Deeply Virtual Compton Scattering (DVCS) experiments conducted at Thomas Jefferson National Accelerator Facility, utilizing Quantum Deep Neural Networks (QDNNs). The analysis employs the standard Belitsky, Kirchner, and Müller formalism at twist-two, complemented by a fitting procedure designed to minimize model dependence in a manner analogous to conventional local fits. A pseudodata extraction test of the CFFs is performed using both Classical Deep Neural Networks (CDNNs) and QDNNs, with a detailed comparative analysis. Results indicate that QDNNs can outperform CDNNs in particular cases, offering enhanced predictive accuracy and precision even with limited model complexity. Motivated by this, we develop a metric to quantify the extent of the quantum advantage based on characteristics of DVCS experimental data. These findings underscore the promising role of QDNNs in advancing future investigations into multidimensional parton distributions and hadronic physics.
@@ I Introduction
Exploring the internal structure of the nucleon has been a longstanding goal in hadronic physics, with deeply virtual exclusive processes providing a crucial window into its partonic composition. Among these, Deeply Virtual Compton Scattering (DVCS) stands out as a uniquely sensitive probe of Generalized Parton Distributions (GPDs), which encode the spatial and momentum distributions of quarks and gluons. By linking longitudinal momentum fractions to transverse spatial coordinates, GPDs provide critical insights into the orbital angular momentum carried by partons, bridging the gap between experimental observables and the deeper theoretical understanding of nucleon structure.
@@
Despite their fundamental importance, GPDs cannot be directly extracted from experimental measurements due to their integral appearance in DVCS amplitudes. Instead, Compton Form Factors (CFFs) emerge as a more experimentally accessible alternative, encapsulating the convolution of GPDs with perturbative coefficient functions. These complex-valued functions serve as indirect yet powerful observables, enabling the study of quark and gluon distributions through measurable quantities such as cross sections and asymmetries. This may necessitate the advancement of analysis techniques to go beyond classical methods, leveraging approaches capable of capturing the quantum correlations inherent in the DVCS process. In this context, quantum deep neural networks (QDNNs) offer a promising approach to the systematic extraction of quantum information pertaining to partonic degrees of freedom, enhancing our ability to measure and interpret the internal quantum structure of the nucleon.
@@
QDNNs are neural network architectures that take advantage of quantum mechanics to process information. Unlike classical deep learning models that rely on traditional computation, QDNNs employ quantum circuits as trainable components, allowing them to explore complex, high-dimensional function spaces more efficiently. Libraries such as PennyLane and Quantum TensorFlow provide frameworks to construct and train QDNNs using hybrid quantum-classical approaches. These frameworks enable the encoding of classical or quantum data into quantum states, allowing manipulation through quantum circuits, and extraction of results via measurement processes. PennyLane is designed to integrate quantum computing with machine learning and automatic differentiation techniques, allowing users to build quantum nodes (or “QNodes”) that can be optimized in conjunction with classical deep learning architectures. Quantum TensorFlow extends Google’s TensorFlow to incorporate quantum computation, enabling quantum variational circuits to be trained using classical gradient descent methods. These tools provide a seamless way to incorporate quantum algorithms into machine learning tasks, facilitating the study of quantum-enhanced feature extraction, pattern recognition, and regression tasks, particularly those of a quantum nature.
@@
One application of QDNNs is in the extraction of kinematically sensitive parameters, such as Compton Form Factors (CFFs), from helicity amplitudes in deeply virtual Compton scattering (DVCS). In DVCS, the helicity amplitudes encode critical information about the nucleon’s partonic structure, represented as complex-valued functions of kinematic variables like momentum transfer squared $t$, Bjorken scaling variable $x_{B}$, and photon virtuality $Q^{2}$. Extracting CFFs from experimental data is challenging due to limited constraints and the interference with the Bethe-Heitler process, where the final state (electron, nucleon, and photon) arises from two coherent quantum amplitudes that interfere constructively or destructively. This interference, driven by quantum superposition and phase coherence, complicates the separation of DVCS contributions. Traditionally, feed-forward deep neural networks have been used to fit experimental data and extract these parameters. However, classical models may struggle to efficiently capture the complex quantum correlations [1] and interference patterns inherent in the data, especially if the data has large experimental errors. QDNNs offer a novel approach by encoding kinematic inputs into quantum states and processing them with quantum variational circuits. While the quantum mechanical properties of QDNNs, such as entanglement in their circuit architecture, are primarily used as a flexible parameterization for regression, they may provide enhanced expressive power [1] to model the intricate structures in DVCS data compared to classical networks under particular circumstances. This potential advantage warrants further exploration to determine whether QDNNs can better represent the underlying quantum correlations in DVCS processes or simply provide an alternative complementary extraction tool.
@@
By leveraging quantum entanglement and interference, QDNNs can capture nontrivial correlations between observables that are difficult for classical models to learn efficiently. In CFF extraction, classical extraction techniques have to approximate real and imaginary parts separately, possibly missing quantum interference and correlation effects that are intrinsic to the DVCS amplitude. However, a QDNN can process this information holistically by leveraging superposition and quantum state evolution, which naturally preserve these quantum correlations [1]. Additionally, quantum-enhanced feature spaces allow for a more expressive representation of the functional dependencies of helicity amplitudes. This is particularly useful in scenarios where experimental uncertainties and systematic errors obscure direct parameter extraction. Since QDNNs operate in a higher-dimensional Hilbert space, they can encode more complex functional mappings, potentially leading to more accurate determinations of observables like beam-spin and target-spin asymmetries, which are crucial in probing the nucleon structure.
With the rapid advancement of near-term quantum processors, such as those developed by IBM, Google, and Rigetti, experimental physics stands at the threshold of utilizing QDNNs for real-world data analysis. A promising approach involves integrating quantum machine learning with Monte Carlo sampling methods to systematically propagate experimental errors through the quantum model, improving the robustness of extracted parameters like CFFs. The application of Quantum DNNs in libraries like PennyLane and Quantum TensorFlow offers a new paradigm for extracting quantum-mechanical observables in nuclear and high-energy physics. By exploiting quantum feature spaces, entanglement, and parameterized quantum circuits, these models can potentially outperform classical approaches in unraveling the helicity structure of deeply virtual processes like DVCS. As quantum hardware continues to improve, QDNNs may soon become indispensable tools for analyzing quantum field-theoretic processes, leading to deeper insights into the fundamental structure of matter. This approach may enhance the extraction fidelity of critical observables and open new avenues for studying the quantum correlations embedded within the hadronic structure.
@@
In this work, we demonstrate some very basic applications of QDNNs for information extraction and develop some metrics for comparison. Then we perform some extraction testing using a standard local fit that involves fitting a single observable at fixed kinematics, across the independent variable $\phi$, the azimuthal angle between the lepton and hadron scattering planes. Local fitting relies on the chosen framework of general approximation, such as twist truncation, leading-order contributions, and the dominance of some CFFs. At the highest approximation, the parametrization of observables, such as the cross section, in terms of CFFs is not unique as it is convention-dependent since it violates the required symmetries, which are restored by contributions that are formally suppressed by powers of $1/Q$, dubbed higher-twist corrections. In this regard, the extracted CFF depends on the choice of the observable formalism [2, 3, 4, 5, 6, 7].
@@
While local fits offer the advantage of being potentially model-independent, extracting unique CFFs for fixed kinematics remains a formidable challenge, particularly when the number of available observables is limited, reducing the constraints in the fit. Since CFFs can vary freely within a high-dimensional parameter space, standard $\chi^{2}$ minimization methods often suffer from inconsistent local minima, leading to significant uncertainties in their extraction, even in the most rigorous analyses. As experimental datasets continue to grow in both quantity and precision, particularly with the inclusion of multiple polarized observables, the issue of data sparsity is gradually improving. However, to fully capitalize on these advances, it is equally crucial to refine and enhance extraction techniques, ensuring that they are capable of handling the increasing complexity and correlations of the data.
@@
We employ a straightforward approach using simple feedforward architectures and, similar to previous work  [8, 9], incorporate the helicity amplitude formalism directly into the loss function during the backpropagation of error. To account for experimental uncertainties, we propagate errors through a Monte Carlo ensemble of replicas [10], where the cross section data is resampled within its experimental uncertainty. This ensemble generates a distribution of extracted CFFs, with the width of the distribution reflecting both the inherent experimental errors and the uncertainties arising from the fitting procedure.
@@
We focus on the least constrained observable case: the unpolarized DVCS cross section. We adopt the formalism developed by Belitsky and Müller, specifically the BKM10 framework  [4], which defines the structure of the helicity amplitudes used in the extraction. While this approach is tailored to the BKM10 formalism, the extraction method itself is general and can be adapted to alternative theoretical frameworks as needed. Future work will explore an expanded approach aimed at minimizing potential biases introduced by the choice of helicity amplitude formalism. A brief overview of the BKM10 formalism will be provided in the next section.
@@
In this study, we compare the performance of CDNNs and QDNNs in the extraction of CFFs, using a consistent theoretical framework and a new quantum metric. To ensure a direct comparison, we use the BKM10 formalism  [4] for both approaches, maintaining uniformity in the helicity amplitude structure used in the loss function. We also focus on quantifying and matching model complexity. By applying identical pseudo input data and theoretical constraints, we systematically evaluate how QDNNs enhance the extraction process relative to CDNNs, particularly in the context of the various contributions to the extraction error and fit ambiguities. The BKM10 formalism, which will be briefly discussed in the next section, provides a well-established foundation for this comparative study, allowing us to isolate the advantages or disadvantages introduced by quantum machine learning techniques.
The remainder of this article is organized as follows. Section II provides a comprehensive overview of the theoretical framework for Deeply Virtual Compton Scattering (DVCS), detailing the role of Compton Form Factors (CFFs) in the extraction process. Section III introduces Quantum Deep Neural Networks (QDNNs), explaining their architecture and advantages over Classical Deep Neural Networks (CDNNs). Examples are provided that explore quantum classification and regression, comparing QDNN and CDNN performance on basic classification and regression tasks with well-defined metrics using pseudodata. Section IV describes the experimental data from Thomas Jefferson National Accelerator Facility used both to generate pseudodata for testing and as input for the extraction of real experimentally determined CFFs in Sec. VII. Section V details the methodology for CFF extraction, including a comparative analysis of CDNN and QDNN performance using accuracy, precision, and error metrics. Section VI develops a DVCS “quantum qualifier,” a metric designed to predict whether a QDNN will outperform a CDNN in CFF extraction based purely on the characteristics of the experimental DVCS data. Section VII applies the optimized QDNN methodology to real experimental DVCS data, showcasing the practical utility and robustness of the quantum approach in a realistic setting. Finally, Sec. VIII discusses the results and concludes with key findings and future research directions for quantum-enhanced parameter extraction in hadronic physics.
@@ II Theory Framework
To fully appreciate the challenges of extracting CFFs and how quantum models might help, we begin with a review of the theoretical framework that connects GPDs, CFFs, and measurable DVCS observables. The DVCS process studies the virtual Compton scattering in deep-inelastic kinematics at the leading order in perturbative QCD. DVCS is dominated by single-quark scattering, and therefore the amplitude can be expressed in terms of the off-forward parton distributions. In the helicity-independent case, a virtual photon scatters from an unpolarized electron beam of energy $k$ off an unpolarized proton leading to the 4-fold photon electroproduction differential cross section:
@@
The phase space of this process is parameterized by the Bjorken variable, $x_{B}=Q^{2}/2pq$, the four-momentum $q=k-k^{\prime}$ carried by the virtual photon of mass $Q^{2}={-q}^{2}=-(k-k^{\prime})^{2}$, the squared momentum transfer between the initial and final protons $t=\Delta^{2}$ with $\Delta=p^{\prime}-p$ and the lepton energy loss $y=(p\cdot q)/(p\cdot k)$. The azimuthal angle $\phi$ between the leptonic and hadronic planes is defined in the Trento convention [11]. In the cross section, $\alpha=e^{2}/(4\pi)$ is the fine structure constant, and $\epsilon=2x_{B}M/Q^{2}$ where $M$ is the proton mass.
The photon electroproduction cross section is sensitive to the coherent interference of the DVCS amplitude with the Bethe-Heitler (BH) amplitude. The DVCS process, namely, the electroproduction of a photon off a proton with the exchange of a photon with virtuality, $Q^{2}$, much larger than the momentum transfer squared $t$, hypothetically provides one of the cleanest probes of the GPDs. The BH is an easily calculated electromagnetic process, where a real photon is emitted by the initial or the final lepton. Therefore, both DVCS and BH channels contain the same final states, which can not be distinguished experimentally, and the total cross section arises from the modulus squared of the coherent sum of amplitudes:
@@
[…]
At leading order (LO) in perturbative QCD and leading twist (LT) approximation, four independent proton GPDs can be accessed in the DVCS process: $H$, $E$, $\widetilde{H}$ and $\widetilde{E}$. These four GPDs depend on three variables: $x$, $\xi$ and $t$ where the latter two are experimental parameters with the skewness $\xi=x_{B}/(2-x_{B})$ at leading twist and $x\pm\xi$ the light-cone longitudinal momentum fraction of the struck quark before (“$+$” sign) and after (“--” sign) the scattering, such that the difference in the momentum fraction of the quark before absorbing the virtual photon and after the emission of the real photon is given by $2\xi$, resulting in a squared four-momentum transfer $t$ to the nucleon. This gives rise to the division of the GPDs into two distinct regions, each carrying entirely different physical interpretations. In the region where the momentum fraction $x$ exceeds $\xi$ or falls below $-\xi$ the GPDs are interpreted as the amplitudes governing the emission and reabsorption of a parton. Conversely, within the region encompassing $-\xi<x<\xi$, GPDs resemble distribution amplitudes governing the emission or absorption of a parton-antiparton pair.
@@
where $e_{q}$ is the charge of quarks in units of proton charge and the summation runs over all quark flavors $q$. Using the residue theorem, it is possible to decompose the integral (5) so that each CFF contains two real quantities:
@@
[…]
@@
The top sign in equations (5)-(10), corresponds to the unpolarized GPDs ($H,E$) and the bottom sign applies to the polarized GPDs ($\widetilde{H},\widetilde{E}$). Equations (8) and (9) show that observables sensitive to the imaginary part of the CFFs will only contain information along the line $x=\pm\xi$, whereas the real part probes GPD integrals over the momentum fraction $x$. Given the complicated kinematic dependence of the GPDs, retrieving them from CFFs is therefore a major challenge of the field, known as the deconvolution problem [17]. Thus, the maximum model-independent information which can be extracted from the DVCS reaction at leading twist are 8 CFFs, which depend on two variables, $\xi$ and $t$, at QCD leading order. There is an additional $Q^{2}$ dependence in the CFFs (and in the GPDs) if QCD evolution is taken into account [18].
Our extraction uses the unpolarized beam and unpolarized target cross sections relying on the Belitsky-Kirchner-Müller DVCS formulation (BKM10) [4] and we focus on the case where only twist-2 CFFs enter the cross section. In this case, the squared DVCS amplitude is a bilinear combination of the CFFs contained in the coefficient $\mathcal{C}^{DVCS}(\mathcal{F},\mathcal{F}^{*})$:
@@
The double helicity-flip gluonic CFFs amplitudes, $C^{\mathcal{I},n}_{-+}(\mathcal{F_{T}})$, are formally suppressed by $\alpha_{s}$ and will be neglected here. $C^{\mathcal{I},n}_{++}(\mathcal{F})$ and $C^{\mathcal{I},n}_{0+}(\mathcal{F}_{eff})$ are the helicity-conserving and the helicity-changing amplitudes respectively. The helicity-changing amplitudes also contain the contribution from twist-3 CFFs implicit in the effective CFFs ($\mathcal{F}_{eff}$). Thus, we chose a framework where only the helicity-conserving amplitudes are considered, which contain only pure twist-2 CFFs contributions. In this way, we set more constraints on the physics with the same amount of data, giving up to some information at the expense of greater accuracy in the remaining parameters. The helicity-conserving amplitudes are given by:
@@
In this specific scenario, the interference cross section becomes reliant solely on three real-valued CFFs: $\Re e\mathcal{H}$, $\Re e\mathcal{E}$, and $\Re e\widetilde{\mathcal{H}}$. The real parts of the CFFs $\Re e\mathcal{H}$, $\Re e\mathcal{E},$ and $\Re e\widetilde{\mathcal{H}}$ arise from the dispersive integrals of the corresponding GPDs $H$, $E$, and $\widetilde{H}$, representing their convolution with the hard scattering kernel. $\Re e\mathcal{H}$ dominates the DVCS amplitude for unpolarized targets and reflects the spatial-momentum distribution of unpolarized quarks. $\Re e\mathcal{E}$ is sensitive to nucleon helicity-flip and provides access to parton orbital angular momentum via Ji’s sum rule. $\Re e\widetilde{\mathcal{H}}$ encodes the distribution of polarized quarks and governs the longitudinal spin structure, playing a leading role for polarized targets. Additionally, accounting for the constant nature of the pure DVCS cross section in terms of the azimuthal angle, the total number of parameters that need to be extracted is reduced to four. We chose this framework to make the comparison of CDNN and QDNN extractions more direct. The reduction in the parameter count simplifies the extraction process relative to the full eight-parameter scenario, while still preserving the most relevant aspects of the phenomenological information. This simplified parameter space facilitates a more manageable data analysis and modeling, making a more straightforward demonstration of the extraction steps of the essential parameters from experimental data. A separate future paper will investigate directly the quantum advantage in extracting the complex-valued CFFs.